# A Competition Winning Deep Reinforcement Learning Agent in microRTS

### v1

@@ 3.3 Transfer learning
We trained the squnet models with fewer steps because of time constraints and skipped the predominantly cost-based reward phase because it didn’t help train the base model (Table 13). The 2 models (trained on maps of up to size 32x32 and 64x64) managed only a 40% win rate, never beating CoacAI or Mayari. These models were policies of last resort.
@@
3.5 Behavior cloning bootstrapped training
In follow-up work after the RAISocketAI submission, we wanted to train a model that (1) didn’t require the shaped rewards and reward scheduling, (2) could be trained in fewer steps and less time, and (3) could defeat prior competition winners on the largest maps. We opted for a neural architecture between DoubleCone and squnet: a nested downscaling residual block each of stride 4, so the bottom block scales the input down to 1/16th the original size. At each downscaling level, there were multiple residual blocks (6 at full resolution, split evenly by the downscaling block; 4 at 1/4 resolution, split evenly by the 1/16 downscaling block; and 4 at 1/16 resolution) (Table 15). This architecture theoretically has a 128x128 receptive field while using 25% fewer operations than DoubleCone at inference time. On the largest Open map ((4)BloodBath.scmB), this is 6 times more computations than squnet-64. Therefore, this neural architecture won’t be usable in a competition given the same hardware constraints as the 2023 competition.
@@ 3.5 Behavior cloning bootstrapped training
Next, we tried imitation learning to bootstrap the model, similar to Vinyals et al. (2019). We got rid of the three rewards, opting for only the win-loss reward. microRTS doesn’t have human replays, so we used playthroughs of the 2021 competition winner Mayari playing against itself, 2020 competition winner CoacAI, and POLightRush (baseline scripted bot and 2017 competition winner) (Table 16). Instead of generating an offline replay dataset, we set the microRTS environment to play bots against each other and these observations and actions were fed into rollouts used for behavior cloning the policy and fitting the value heads:
$L^{BC+VF}(\theta)$ $=\hat{\mathbb{E}}_{t}\left[L^{BC}(\theta)-c_{1}L^{VF}(\theta)\right],$ (4)
$L^{BC}(\theta)$ $=-\frac{1}{\|a_{t}\|}\log\pi_{\theta}(a_{t}|s_{t}),$ (5)
where $c_{1}$ and $L^{VF}$ are the same as in the PPO loss function. The behavior cloning policy loss is the cross-entropy loss between the policy logits and the actions taken by the Mayari bot. We found scaling the loss by the number of units accepting actions allowed the learning rate to be significantly increased. Scaling down the loss by the number of units keeps the losses for all turns roughly similar in scale as otherwise large unit count turns would have much larger losses as each unit’s loss contribution is summed together (each unit’s actions are assumed to be independent).
We trained 3 behavior cloned models (16x16, 32x32, and 64x64) on the same maps for each map size as RAISocketAI training. The 64x64 model used the weights of the 32x32 model as a starting point, while the other two models were randomly initialized. We then used PPO to fine-tune the behavior cloned models on the same maps (Table 17).
4 Results
4.1 Single player round-robin benchmark
Table 1: Single player round-robin benchmark win rates. Win rates above 50% are bolded. (D) DoubleCone model used. (S) squnet model used.
WorkerRush LightRush CoacAI Mayari Overall
@@ 4.1 Single player round-robin benchmark
BWDistantResources32x32
99 (D)
93 (S)
90 (D)
73 (S)
88 (D)
23 (S)
99 (D)
58 (S)
94 (D)
61 (S)
(4)BloodBath.scmB 98 0 0 0 25
AI Average* 99 85 83 87 88
    AI Average uses the DoubleCone (D) results from BWDistantResources32x32.
@@
In a single player round-robin benchmark on the Open maps (Table 1), RAISocketAI beat the competition winners of 2021 (Mayari), 2020 (CoacAI), and 2017 (POLightRush, baseline) on 7 of the 8 maps (winning over 96% of games on these maps). RAISocketAI could only beat the POWorkerRush baseline bot on the largest map, (4)BloodBath.scmB. The DistantResources fine-tuned squnet model performed worse than the DoubleCone model across all opponents, but maintained an over 50% win rate against all but CoacAI. Timeouts didn’t affect results significantly.
4.2 IEEE-CoG 2023 microRTS competition results
The IEEE-CoG 2023 microRTS competition is a round-robin tournament on 12 maps of different sizes and distributions of terrain, resources, and starting units and buildings. 8 Open maps are known beforehand, 4 Hidden maps are only revealed after the competition. The winner is the agent with the highest win rate on the 8 Open maps. Hidden map results are publicly available, but this paper will only discuss the Open maps. For this competition, a total of 11 agents were submitted: 9 programmatic policies, 1 synthesized programmatic policy, and RAISocketAI. The competition also had 6 baselines: (1) RandomBiasedAI (performs actions randomly, biased towards attacking if able), (2) NaiveMCTS (a simple Monte Carlo tree search agent that searches until reaching the time limit) (3) POWorkerRush, (4) POLightRush, (5) 2L (programmatic strategies generated by Moraes et al. (2023), the competition organizers), and (6) the prior competition winner Mayari. The baselines cannot win the competition.
RAISocketAI was declared the winner with the highest win rate (72%) across all submissions (Table 2). RAISocketAI had a higher win rate than all but two baselines: 2L (76%) and Mayari (82%). RAISocketAI had an over 50% win rate versus every opponent including 2L (60%) and Mayari (65%).
Table 2: Win rates of selection of agents in the IEEE-CoG 2023 microRTS competition. Player 1 is the row agent and player 2 is the column agent. Each win rate value is the percentage of games won by player 1. Cells are bolded if the win rate is higher than the opponent’s row win rate. For example, RAISocketAI vs ObiBotKenobi is bolded because 49% is higher than 47%, thus meaning the combined player 1 and 2 win rate is 51% for RAISocketAI vs ObiBotKenobi. Overall includes all agents, including those not shown. Win rates for all agents are shown in Table 26.
@@ 4.2 IEEE-CoG 2023 microRTS competition results
Table 3: RAISocketAI win rates in 2023 competition by opponent and map. Bolded cells are win rates over 50%. Overall includes all agents, including those not shown.
@@
As expected from the single player round-robin benchmark, RAISocketAI does better on smaller maps and dismally on the largest maps (Table 3). However, in the competition, RAISocketAI underperformed against agents already benchmarked in the single player round-robin (14-19% lower win rate against each agent), even accounting for the likely use of the weaker squnet model on BWDistantResources32x32. Breaking down by map, RAISocketAI underperformed against benchmarked agents by 20-40% on 5 maps.
The competition ran jobs, splitting each map into 5 or 10 jobs where each job would run a complete round-robin with all agents on that map playing 2 or 1 games as player 1 and 2 each. For basesWorkers8x8A, on which RAISocketAI underperformed by almost 40%, the competition had 5 jobs. On the first 3 jobs, RAISocketAI won nearly every game. On the last 2 jobs, RAISocketAI lost nearly every game. 1 or 2 jobs per underperforming map appear to have outlier low win rates for RAISocketAI (Table 27).
4.3 Behavior cloning results
We created 2 additional agents: (1) only behavior cloning (RAI-BC) and (2) followed by PPO fine-tuning (RAI-BC-PPO). Each agent consisted of the models trained on their respective map sizes (16x16, 32x32, and 64x64). These agents do not have any map-specific models.
Table 4: RAI-BC-PPO win rate in a single player round-robin benchmark. Win rates above 50% are bolded.
POWorkerRush POLightRush CoacAI Mayari Overall
@@ 4.3 Behavior cloning results
RAI-BC had a 71% win rate, doing well against the POLightRush baseline (96%, better than RAISocketAI) and respectably against Mayari (44%) (Table 28). On the largest map, RAI-BC manages to occasionally beat POLightRush (63%), CoacAI (20%), and Mayari (40%) compared to RAISocketAI’s 0% against these 3 more difficult opponents.
Once fine-tuned with PPO, RAI-BC-PPO obtains an RAISocketAI-comparable 88% win rate (Table 4). RAI-BC-PPO generally improves upon RAI-BC’s win rates on each map and against each opponent. However, the biggest exceptions are POLightRush on TwoBasesBarracks16x16 (from 100% to 0%) and the largest map where the fine-tuned model can no longer beat CoacAI and Mayari.
@@
5 Discussion
5.1 Improving inference time in microRTS competitions
@@
Improving inference time is critical to matching benchmark results in a competition. We suggest 3 improvements (2 for agents, and 1 for the competition organizers): (1) use fast inference runtime providers like OpenVINO for ONNX Runtime, (2) continue to train agents using the smaller squnet models (possibly with behavior cloning to bootstrap training), and (3) replace the fixed per turn timeout tolerance in the competition with an overtime budget. For DoubleCone, Limburg (2023) found using OpenVINO could have made inference 2-3 times faster in the LUX competition. This would likely make running DoubleCone or the larger squnet models on all maps except the largest feasible for the competition. An overtime budget for an entire match instead of the 20 ms per turn tolerance will help agents deal with environment instabilities. For example, RAI-BC-PPO timed out in 11% of games on (4)BloodBath.scmB, despite averaging 55 ms/turn and going over 100 ms in only 0.016% of turns (averaging less than 1 over 100 ms turn per game). An overage budget of even 1 second per game would likely prevent most timeouts.
5.2 Training on larger maps
@@
We hoped imitation learning would mitigate these issues by providing a model that generates non-zero win-loss rewards and reasonable observation-action pairs. However, during PPO fine-tuning, a training policy that initially won 40-50% of training games, dropped to 20% midway through training. It eventually recovers to winning 40% of training games; however, the fine-tuned policy had a worse evaluation win rate than the initial supervised policy. This training curve differs from the smaller map fine-tuning where the training policy quickly won 60% of training games and improved upon the evaluation win rate.
PPO (and DRL algorithms in general) have many hyperparameters that need to be tuned before a model is able to learn, and our hyperparameter search is ongoing. Other algorithms might perform better from a behavior cloned start. For example, Advantage Actor Critic (A2C) (Mnih et al., 2016) uses a loss function similar to behavior cloning, and thus policy training might more smoothly transition from behavior cloning to DRL.
Extending the curriculum to include more game states could also improve large map training. For example, varied existing agents can be used to advance the game before switching to a training agent to finish the game (Uchendu et al., 2023). Another way is to utilize a prioritized fictitious self-play mechanism used by Vinyals et al. (2019) that prioritizes training the agent on the most difficult prior agent checkpoints.
5.3 Behavior cloning, transfer learning, and academic competitions
@@
There are several ways to make DRL more feasible in a competition and educational setting: (1) focus on smaller maps, (2) fine-tune pretrained models from DRL or behavior cloning, (3) transfer an existing model to new maps, or (4) use a significantly smaller neural architecture. The largest map took 19 GPU-days to train for RAISocketAI, 15 GPU-days for RAI-BC, and 34 GPU-days for RAI-BC-PPO. Over two-thirds of training time for RAI-BC-PPO was spent training on the largest map to little benefit. Huang et al. (2021) trained an agent for player 1 on a single 16x16 map in 2.5 days. RAI-BC-PPO trained on the 5 Open maps up to size 16x16 in about 7.5 days.
@@
RAISocketAI’s DoubleCone and RAI-BC-PPO’s deep squnet are relatively large neural networks, each at around 5 million parameters. Huang et al. (2021)’s best performing policies each used fewer than 1 million parameters. These smaller networks are quicker to train and have faster inference time, and there’s no definitive evidence so far that they are worse than larger networks.
6 Conclusion
RAISocketAI is the first DRL agent to win a microRTS competition. It demonstrates that an iterative training process of fine-tuning and transfer learning is effective for creating competitive DRL agents. Such a training process can be used by resource-constrained researchers and students to create novel DRL agents for future competitions and experiments. Fine-tuning behavior cloning with PPO is a promising way to create competitive DRL agents without needing to handcraft shaped reward functions.
7 Reproducibility Statement
Our Methods section describes the architectures and training processes for RAISocketAI, RAI-BC, and RAI-BC-PPO. We give additional details in Appendices: Appendix B neural network architectures, Appendix C initial training details, Appendix D transfer learning details, Appendix E squnet learning details, and Appendix F behavior cloning details. Appendix H describes the setup of the single player round-robin tournaments for RAISocketAI (Section 4.1), RAI-BC, and RAI-BC-PPO (Section 4.3). The microRTS competition requires agents to be open sourced. Our open-sourced code repository includes a link to the archive used for the competition and instructions on how to run the submission in the competition environment [^5] https://github.com/sgoodfriend/rl-algo-impls/tree/main/rl_algo_impls/microrts.

### v2

@@ III-D Squnet Training
We trained the squnet models with fewer steps because of time constraints and mostly didn’t use the cost-based reward because it didn’t help train the base model. The 2 models (trained on maps of up to size 32x32 and 64x64, respectively) managed only a 40% win rate, never beating CoacAI or Mayari. These models were policies of last resort.
@@
III-E Behavior Cloning Bootstrapped Training
In follow-up work after the RAISocketAI submission, we wanted to train a model that (1) didn’t require the shaped rewards and reward scheduling, (2) could be trained in fewer steps and less time, and (3) could defeat prior competition winners on the largest maps. We opted for a neural architecture between DoubleCone and squnet: a nested downscaling residual block each of stride 4, so the bottom block scales the input down to 1/16th the original size. At each downscaling level, there were multiple residual blocks (6 at full resolution, split evenly by the downscaling block; 4 at 1/4 resolution, split evenly by the 1/16 downscaling block; and 4 at 1/16 resolution). This architecture theoretically has a 128x128 receptive field while using 25% fewer operations than DoubleCone at inference time. On the largest Open map ((4)BloodBath.scmB), this is 6 times more computations than squnet-64. Therefore, this neural architecture wouldn’t be usable in a competition given the same hardware constraints as the 2023 competition.
@@ III-E Behavior Cloning Bootstrapped Training
Next, we tried imitation learning to bootstrap the model, similar to [3]. We got rid of the three rewards, opting for only the win-loss reward. microRTS doesn’t have human replays, so we used playthroughs of the 2021 competition winner Mayari playing against itself, 2020 competition winner CoacAI, and POLightRush (baseline scripted bot and 2017 competition winner). Instead of generating an offline replay dataset, we set the microRTS environment to play bots against each other and these observations and actions were fed into rollouts used for behavior cloning the policy and fitting the value heads:
$L^{\text{BC+VF}}(\theta)$ $=\hat{\mathbb{E}}_{t}\left[L^{\text{BC}}(\theta)+c_{1}L^{\text{VF}}(\theta)\right],$ (4)
$L^{\text{BC}}(\theta)$ $=-\frac{1}{\|a_{t}\|}\log\pi_{\theta}(a_{t}|s_{t}),$ (5)
where $c_{1}$ and $L^{\text{VF}}$ are the same as in the PPO loss function. The behavior cloning policy loss is the cross-entropy loss between the policy logits and the actions taken by the Mayari bot. We found scaling the loss by the number of units accepting actions allowed the learning rate to be significantly increased. Scaling down the loss by the number of units keeps the losses for all turns roughly similar in scale as otherwise large unit count turns would have much larger losses as each unit’s loss contribution is summed together (each unit’s actions are assumed to be independent).
We trained 3 behavior cloned models (16x16, 32x32, and 64x64) on the same maps for each map size as RAISocketAI training. The 64x64 model used the weights of the 32x32 model as a starting point, while the other two models were randomly initialized. We then used PPO to fine-tune the behavior cloned models on the same maps.
IV Results
IV-A Single Player Round-robin Benchmark
In a single player round-robin benchmark on the Open maps (Table I), RAISocketAI beat the competition winners of 2021 (Mayari), 2020 (CoacAI), and 2017 (POLightRush, baseline) on 7 of the 8 maps (winning over 96% of games on these maps). RAISocketAI could only beat the POWorkerRush baseline bot on the largest map, (4)BloodBath.scmB. The DistantResources fine-tuned squnet model performed worse than the DoubleCone model across all opponents, but maintained an over 50% win rate against all but CoacAI. Timeouts didn’t affect results significantly.
TABLE I: Single player round-robin benchmark win rates. Win rates over 50% are bolded. Higher win rates are redder. Lower win rates are bluer.
WorkerRush
LightRush
CoacAI
Mayari
Overall
@@ IV-A Single Player Round-robin Benchmark
BWDistantResources32x32a  99  /93  90  /73  88  /23  99  /58  94  b /61
(4)BloodBath.scmB  98   0   0   0   25  c
AI Averaged  99   85   83   87   88
  • a
    Two models were trained for BWDistantResources32x32: DoubleCone (1st number) and squnet (2nd number). Cell color based on DoubleCone results.
  • b
@@
  • d
    AI Average uses the DoubleCone (1st number) results from BWDistantResources32x32.
IV-B IEEE-CoG 2023 microRTS Competition Results
The IEEE-CoG 2023 microRTS competition is a round-robin tournament on 12 maps of different sizes and distributions of terrain, resources, and starting units and buildings. 8 Open maps are known beforehand, 4 Hidden maps are only revealed after the competition. The winner is the agent with the highest win rate on the 8 Open maps. Hidden map results are publicly available, but this paper will only discuss the Open maps. For this competition, a total of 11 agents were submitted: 9 programmatic policies, 1 synthesized programmatic policy, and RAISocketAI. The competition also had 6 baselines: (1) RandomBiasedAI (performs actions randomly, biased towards attacking if able), (2) NaiveMCTS (a simple Monte Carlo tree search agent that searches until reaching the time limit) (3) POWorkerRush, (4) POLightRush, (5) 2L (programmatic strategies generated by [16], the competition organizers), and (6) the prior competition winner Mayari. The baselines cannot win the competition.
RAISocketAI was declared the winner with the highest win rate (72%) across all submissions (Table II). RAISocketAI had a higher win rate than all but two baselines: 2L (76%) and Mayari (82%). RAISocketAI had an over 50% win rate versus every opponent including 2L (60%) and Mayari (65%).
TABLE II: Win rates of selection of agents in the IEEE-CoG 2023 microRTS competition. The row agent is player 1, while the column agent is player 2. The win rate value is the percentage of games won by player 1. Cells are bolded if the win rate is higher than the opponent’s row win rate.
@@ IV-B IEEE-CoG 2023 microRTS Competition Results
  • a
    Overall includes all agents, including those not shown.
  • b
    RAISocketAI vs ObiBotKenobi is bolded because 49% is higher than 47%, thus meaning the combined player 1 and 2 win rate is 51% for RAISocketAI vs ObiBotKenobi.
As expected from the single player round-robin benchmark, RAISocketAI does better on smaller maps and dismally on the largest maps (Table III). However, in the competition, RAISocketAI underperformed against agents already benchmarked in the single player round-robin (14-19% lower win rate against each agent), even accounting for the likely use of the weaker squnet model on BWDistantResources32x32. Breaking down by map, RAISocketAI underperformed against benchmarked agents by 20-40% on 5 maps.
TABLE III: RAISocketAI win rates in 2023 competition by opponent and map. Win rates over 50% are bolded. Higher win rates are redder. Lower win rates are bluer.
@@
  • a
    Overall includes all agents, including those not shown.
The competition ran jobs splitting each map into 5 or 10 jobs where each job would run a complete round-robin with all agents on that map playing 2 or 1 games, respectively, as player 1 and 2 each. For basesWorkers8x8A, on which RAISocketAI underperformed by almost 40%, the competition had 5 jobs. On the first 3 jobs, RAISocketAI won nearly every game. On the last 2 jobs, RAISocketAI lost nearly every game. 1 or 2 jobs per underperforming map appear to have outlier low win rates for RAISocketAI (Table IV).
TABLE IV: RAISocketAI win rates split by competition job. Outlier jobs are bolded.
basesWorkers8x8A
FourBasesWorkers8x8
NoWhereToRun9x8
basesWorkers16x16A
TwoBasesBarracks16x16
DoubleGame24x24
BWDistantResources32x32
(4)BloodBath.scmB
Overall
 98   97   97   100   94   99   49   34
 100   94   97   100   4   96   53   38
 100   95   33   100   92   100   54   38
 5   92   94   98   94   97   58   39
 28   95   100   100   94   9   58   19
 100   39
 97   19
 30   44
 97   36
 95   36
Averagea  66   95   84   100   75   80   54   34   74
No Outliersb  99   95   97   100   93   98   54   38   84
  • a
    Average is the average win rate for all jobs.
  • b
    No Outliers is the average win rate for all jobs excluding "outlier" jobs.
IV-C Behavior Cloning Results
We created 2 additional agents: (1) only behavior cloning (RAI-BC) and (2) a PPO fine-tuned follow-up of RAI-BC (RAI-BC-PPO). Each agent consisted of the models trained on their respective map sizes (16x16, 32x32, and 64x64). These agents do not have any map-specific models.
RAI-BC had a 71% win rate, doing well against the POLightRush baseline (96%, better than RAISocketAI) and respectably against Mayari (44%). On the largest map, RAI-BC manages to occasionally beat POLightRush (63%), CoacAI (20%), and Mayari (40%) compared to RAISocketAI’s 0% against these 3 more difficult opponents.
Once fine-tuned with PPO, RAI-BC-PPO obtains an RAISocketAI-comparable 88% win rate (Table V). RAI-BC-PPO generally improves upon RAI-BC’s win rates on each map and against each opponent. However, the biggest exceptions are POLightRush on TwoBasesBarracks16x16 (from 100% to 0%) and the largest map where the fine-tuned model can no longer beat CoacAI and Mayari.
TABLE V: RAI-BC-PPO win rate in a single player round-robin benchmark. Win rates over 50% are bolded. Higher win rates are redder. Lower win rates are bluer.
POWorkerRush
POLightRush
CoacAI
Mayari
Overall
@@ IV-C Behavior Cloning Results
V Discussion
V-A Improving Inference Time in microRTS Competitions
@@
Improving inference time is critical to matching benchmark results in a competition. We suggest 3 improvements (2 for agents, and 1 for the competition organizers): (1) use fast inference runtime providers like OpenVINO for ONNX Runtime, (2) train using smaller models (possibly using behavior cloning or policy distillation from larger trained models to bootstrap training), and (3) replace the fixed per turn timeout tolerance in the competition with an overtime budget. For DoubleCone, [11] found using OpenVINO could have made inference 2-3 times faster in the LUX competition. This would likely make running DoubleCone and the larger squnet models on all but the largest maps feasible for the competition. An overtime budget for an entire match instead of the 20 ms per turn tolerance will help agents deal with environment instabilities. For example, RAI-BC-PPO timed out in 11% of games on (4)BloodBath.scmB, despite averaging 55 ms/turn and going over 100 ms in only 0.016% of turns (averaging less than 1 over 100 ms turn per game). An overage budget of even 1 second per game would likely prevent most timeouts.
V-B Training on Larger Maps
@@
We hoped imitation learning would mitigate these issues by providing a model that generates non-zero win-loss rewards and reasonable observation-action pairs. However, during PPO fine-tuning, a training policy that initially won 40-50% of training games, dropped to 20% midway through training. It eventually recovers to winning 40% of training games; however, the fine-tuned policy had a worse evaluation win rate than the initial supervised policy. This training curve differs from the smaller map fine-tuning where the training policy quickly won 60% of training games and improved upon the evaluation win rate. When transitioning to larger maps, [11] used a teacher KL-divergence loss term to keep the policy from diverging too far from the prior trained policy. Using a similar teacher loss term when fine-tuning the behavior cloned policy could help the policy maintain its performance.
Extending the curriculum to include more game states could also improve large map training. For example, varied existing agents can be used to advance the game before switching to a training agent to finish the game [17]. Another way is to utilize a prioritized fictitious self-play mechanism used by [3] that prioritizes training the agent on the most difficult prior agent checkpoints.
While the squnet neural network architectures theoretically have a large enough receptive field to handle the largest map, using a smaller-map trained model as a starting point for transfer learning to the largest maps was not effective. One possible explanation is that the model learns behaviors on a smaller size scale and is not able to generalize behaviors such as resource gathering or enemy pursuit to larger maps. If units were instead processed as entities in a transformer architecture, the agent could process relationships between entities across the map. AlphaStar [3] used a transformer entity encoder as part of its processing.
V-C Academic Competition and Research Considerations
@@
There are several ways to make DRL more feasible in a competition and educational setting: (1) focus on smaller maps, (2) fine-tune pretrained models from DRL or behavior cloning, (3) transfer an existing model to new maps, or (4) use a significantly smaller neural network architecture. The largest map took 19 GPU-days to train for RAISocketAI, 15 GPU-days for RAI-BC, and 34 GPU-days for RAI-BC-PPO. Over two-thirds of training time for RAI-BC-PPO was spent training on the largest map to little benefit. Reference [6] trained an agent for player 1 on a single 16x16 map in 2.5 days. Training RAI-BC followed by RAI-BC-PPO on the 5 Open maps up to size 16x16 took about 7.5 days.
@@
RAISocketAI’s DoubleCone and RAI-BC-PPO’s deep squnet are relatively large neural networks, each at around 5 million parameters. Reference [6]’s best performing policies each used fewer than 1 million parameters. These smaller networks are quicker to train and have faster inference time, and there’s no definitive evidence so far that they are worse than larger networks. Reference [18] uses an entity-based transformer exclusively and demonstrates it can train an agent to beat CoacAI on a 10x10 map in 20 minutes using a single consumer-grade GPU.
VI Conclusion
RAISocketAI is the first DRL agent to win the microRTS competition in 6 iterations. It demonstrates that an iterative training process of fine-tuning and transfer learning is effective for creating competitive DRL agents. Such a training process can be used by resource-constrained researchers and students to create novel DRL agents for future competitions and experiments. Fine-tuning behavior cloning with PPO is a promising way to create competitive DRL agents without needing to handcraft shaped reward functions.
Acknowledgments
We thank the IEEE-CoG 2023 microRTS competition organizers Rubens O. Moraes and Levi H. S. Lelis for their work running the competition, helping us work through performance concerns in the lead-up to our submission, and providing suggestions for this paper. We thank Chang Liu for reading and providing feedback on drafts of this paper.